OpenStudy (anonymous):

Find dy/dx for y = ln (3x^2-4x)

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(f(x))]=\frac{f'(x)}{f(x)}\] in your case \(f(x)=3x^2-4x\) and \(f'(x)=6x-4\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw